### v2

# Improved Space Bounds for Subset Sum

Abstract
More than 40 years ago, Schroeppel and Shamir presented an algorithm that solves the Subset Sum problem for $n$ integers in time $O^{*}(2^{0.5n})$ and space $O^{*}(2^{0.25n})$. The time upper bound remains unbeaten, but the space upper bound has been improved to $O^{*}(2^{0.249999n})$ in a recent breakthrough paper by Nederlof and Węgrzycki (STOC 2021). Their algorithm is a clever combination of a number of previously known techniques with a new reduction and a new algorithm for the Orthogonal Vectors problem.
In this paper, we improve the space bound by Nederlof and Węgrzycki to $O^{*}(2^{0.246n})$ and also simplify their algorithm and its analysis. We achieve this by using an idea, due to Howgrave-Graham and Joux, of using a random prime number to filter the family of subsets. We incorporate it into the algorithm by Schroeppel and Shamir and then use this amalgam inside the representation technique. This allows us to reduce an instance of Subset Sum to a larger number of instances of weighted orthogonal vector.
1 Overview
In this paper, we study the well-known Subset Sum problem and its parameterized version, $k$-SUM. In Subset Sum, given a set of $n$ integers and a target integer $t$ the goal is to check whether there is a subset of them that sum up to $t$. It is common to assume that the absolute value of all input integers is at most $2^{O(n)}$ (one can achieve this using hashing techniques).
In the $k$-SUM problem, the goal is to check whether some $k$ of input integers sum to zero. We consider a slightly different formulation of $k$-SUM: given $k$ sequences $A_{1},\dotsc,A_{k}$, each containing $n$ integers of absolute value at most $n^{O(1)}$, check whether there exist indices $i_{1},\dotsc,i_{k}$ such that [^1] These two versions of $k$-SUM are reduced to each other. One direction is easy: take $k$ copies of the array $A$. For the other direction, one can use the color coding technique [3]: for every element $A$, add it to $A_{i}$ for random $i\in[k]$; then, if a solution for the original problem exists, it survives in the resulting instance with constant probability. This randomized reduction can also be derandomized [3]. $A_{1}[i_{1}]+\dotsb+A_{k}[i_{k}]=0$. It is common to further assume that the bit-length of all integers in the $k$-SUM problem is at most $k\log n+O(\log\log n)$. This can be achieved by the standard fingerprinting technique.
1.1 Known Results
1.1.1 Time Complexity
$2$-SUM can be solved in time $\widetilde{O}(n)$ using binary search (the $\widetilde{O}(\cdot)$ notation hides polylogarithmic factors): sort $A_{1}$; then, for each $1\leq i_{2}\leq n$, check whether $A_{1}$ contains an element $-A_{2}[i_{2}]$. Another well known algorithm is based on the two pointers method that proceeds as follows. Sort $A_{1}$ and $A_{2}$ and let $i_{1}=1$ and $i_{2}=n$. Then, keep repeating the following: if $i_{1}>n$ or $i_{2}<1$, stop; if $A_{1}[i_{1}]+A_{2}[i_{2}]=0$, return $(i_{1},i_{2})$; if $A_{1}[i_{1}]+A_{2}[i_{2}]>0$, decrement $i_{2}$ (as $A_{2}[i_{2}]$ is too large: its sum with the currently smallest element of $A_{1}$ is positive); if $A_{1}[i_{1}]+A_{2}[i_{2}]<0$, increment $i_{1}$ (as $A_{1}[i_{1}]$ is too small: its sum with the currently largest element of $A_{2}$ is negative).
The algorithm for $2$-SUM allows to solve $k$-SUM in time $\widetilde{O}(n^{\lceil k/2\rceil})$ via the following reduction. Given $k$ arrays $A_{1},\dotsc,A_{k}$, split them into two halves: $A_{1},\dotsc,A_{\lfloor k/2\rfloor}$ and $A_{\lceil k/2\rceil},\dotsc,A_{k}$. Populate arrays $B_{1}$ and $B_{2}$ with all sums of elements from the first and second halves, respectively. Then, it remains to solve $2$-SUM for $B_{1}$ and $B_{2}$.
For $3$-SUM, Chan [11] proved a slightly better upper bound $O(n^{2}\log\log^{O(1)}n/\log^{2}n)$. It is a major open problem whether $3$-SUM can be solved in time $O(n^{2-\varepsilon})$ and the $3$-SUM hypothesis, stating that this is impossible, is a popular hypothesis in the field of fine-grained complexity. It is also currently unknown whether $k$-SUM can be solved in time $\widetilde{O}(n^{\lceil k/2\rceil-\varepsilon})$ for any $k\geq 3$ and $\varepsilon>0$.
The $2$-SUM algorithm described above is also at the core of the strongest known upper bound $O^{*}(2^{n/2})$ for Subset Sum ($O^{*}(\cdot)$ hides polynomial factors) presented by Horowitz and Sahni [15] almost fifty years ago. To get this running time, partition $A$ into two halves of size $n/2$; then, populate arrays $A_{1}$ and $A_{2}$ (of size $2^{n/2}$) with sums of all subsets of the two halves, respectively; then, it remains to solve $2$-SUM for $A_{1}$ and $A_{2}$. It is an important open problem to solve Subset Sum in $O^{*}(2^{(1/2-\varepsilon)n})$ for a constant $\varepsilon>0$. Nevertheless, Subset Sum can be solved in time $2^{n/2}/\operatorname{poly}(n)$ [13].
1.1.2 Space Complexity
The algorithms that solve $k$-SUM and Subset Sum via a reduction to $2$-SUM have high space complexity: for $k$-SUM, it is $O(n^{\lceil k/2\rceil})$, whereas for Subset Sum it is $O(2^{n/2})$. It is natural to ask whether one can lower the space complexity or whether it is possible to trade off time and space. Lincoln et al. [19] gave a positive answer for $k$-SUM: one can solve $3$-SUM in time $O(n^{2})$ and space $O(\sqrt{n})$ as well as $k$-SUM (for $k\geq 4$) in time $O(n^{k-2+2/k})$ and space $O(\sqrt{n})$. For Subset Sum, the well-known algorithm by Schroeppel and Shamir [25] solves it in time $O^{*}(2^{n/2})$ and space $O(2^{n/4})$: they reduce Subset Sum to $4$-SUM and note that $4$-SUM can be solved in time $\widetilde{O}(n^{2})$ and space $O(n)$ (since all pairwise sums of two sorted sequences of length $n$ can be enumerated in time $\widetilde{O}(n^{2})$ and space $O(n)$ using a priority queue data structure). Just recently, Nederlof and Wegrzycki improved the space complexity to $O(2^{0.249999n})$ [23].
1.1.3 Proof Complexity
It is easy to certify a yes-instance of $k$-SUM or Subset Sum: just write down a solution; it can be checked in deterministic time $O(n)$. Certifying no-instances is trickier. Carmosino et al. [10] showed that, for $3$-SUM, there are proofs of size $\widetilde{O}(n^{1.5})$ that can be deterministically verified in time $\widetilde{O}(n^{1.5})$. By allowing the verifier to be probabilistic, Akmal et al. [2] presented a proof system for $3$-SUM where the proofs size and the verification time is $\widetilde{O}(n)$. For $k$-SUM, they give an upper bound $\widetilde{O}(n^{k/3})$. The corresponding proof system is known as Merlin–Arthur protocol: a computationally unbounded prover (Merlin) prepares a proof and sends it to a probabilistic verifier (Arthur) who needs to be able to check the proof quickly with small constant error probability. We say that a problem can be solved in Merlin–Arthur time $T(n)$, if there exists a protocol where both the proof size and the verification time do not exceed $T(n)$.
For Subset Sum, Nederlof [22] proved an upper bound $O^{*}(2^{(1/2-\varepsilon)n})$, for some $\varepsilon>10^{-3}$, on Merlin–Arthur time. Akmal et al. [2] improved the bound to $O^{*}(2^{n/3})$.
1.1.4 Cryptographic Primitives
In cryptography, the Subset Sum problem, with its simple description and presumed intractability, has been considered as an alternative to more traditional challenging problems like factoring and discrete logarithm. Starting with the introduction of the Merkle–Hellman [21] cryptosystem in 1978, numerous encryption methods inspired by Subset Sum have been proposed. In 1990, Odlyzko [24] showed how to compromise them. In 1996, Impagliazzo and Naor [17] presented provably-secure primitives based on average-case hardness of Subset Sum. In 1997, Ajtai and Dwork [1] developed a provably-secure cryptosystem based on worst-case hardness of lattice problems that is inspired by Subset Sum. A later study by Lyubashevsky et al. [20] presents a secure encryption method directly based on average-case hardness of Subset Sum.
As cryptographic primitives described above are based on the average-case hardness of Subset Sum, it is natural to ask how fast one can solve random instances of the problem. In 2010, Howgrave-Graham and Joux [16] introduced an algorithm solving Subset Sum on random instances in $O^{*}(2^{0.337n})$ time. A year later, that result was improved to $O^{*}(2^{0.291n})$ by Becker et al. [8], and in 2020, Bonnetain et al. [9] further refined it to $O^{*}(2^{0.283n})$. Inspired by the representation technique introduced in [16], a number of time-space tradeoffs for Subset Sum were established [16, 14, 8, 4]. Later, that technique was used to improve worst-case algorithms for Subset Sum for a class of instances with bounded maximum number of subsets with the same sum [5, 6]. Also, in 2017, Bansal et al. [7] presented an algorithm with running time $O^{*}(2^{0.86n})$ and polynomial space in the model where an algorithm is given access to an exponentially long sequence of random bits. In 2022, the exponential random bits assumption was removed by Chen et al. [12].
1.2 New Results
The main result of the paper is an improved space upper bound for Subset Sum.
Theorem 1.1.
There exists a Monte Carlo algorithm with constant success probability that solves Subset Sum for instances with $n$ integers of absolute value at most $2^{O(n)}$ in time $O^{*}(2^{0.5n})$ and space $O^{*}(2^{0.246n})$.
We achieve this by using an idea, due to Howgrave-Graham and Joux [16], of using a random prime number to filter the family of subsets. This allows us to improve and simplify the the algorithm by Nederlof and Węgrzycki [23]. Below, we give a brief overview of the proof.
Howgrave-Graham and Joux [16] present the following idea of guessing a part of a solution modulo a random prime number. We demonstrate their idea as an Arthur–Merlin protocol. In such protocols, the randomness is shared (also known as public coins) and the verifier is deterministic: upon receiving the verifier’s randomness, the prover prepares a proof and sends it to the verifier who checks it deterministically with small error probability (taken over public randomness). To find integers $i_{1},i_{2},i_{3},i_{4}$ such that $A_{1}[i_{1}]+A_{2}[i_{2}]+A_{3}[i_{3}]+A_{4}[i_{4}]=0$, generate a random prime $p\leq n$ and guess $(A_{1}[i_{1}]+A_{2}[i_{2}])\bmod p$. This gives an Arthur–Merlin protocol with proof size $\log_{2}n$ and verification time $\widetilde{O}(n)$. It has a number of interesting consequences:
  1. 1.
    Arthur–Merlin protocols are efficiently parallelizable: one just enumerates possible proofs on different processors.
  2. 2.
    Using the standard reduction from Subset Sum to 4-SUM, one gets an Arthur–Merlin protocol for Subset Sum with proof size $n/4$ and verification time $O^{*}(2^{n/4})$.
  3. 3.
    By enumerating all proofs of size $n/4$ in time $O^{*}(2^{n/4})$ and verifying each of them in time $O^{*}(2^{n/4})$, one gets an algorithm for Subset Sum with the same upper bounds on time and space as in the algorithm by Schroeppel and Shamir. At the same time, the new algorithm does not need to use the priority queue data structure.
The algorithm by Nederlof and Węgrzycki proceeds roughly as follows. Let $I\subseteq\mathbb{Z}$, $|I|=n$, be an instance of Subset Sum. Assume that it is a yes-instance and that a solution $S\subseteq I$ has size $n/2$. Assume further that there exists a subset $M\subseteq I$ of size $\Theta(n)$ such that $|S\cap M|=|M|/2$ and $M$ is a perfect mixer: the weights of all subsets of $M$ are (pairwise) distinct. Partition $I\setminus M$ into two parts $L$ and $R$ of equal size, hence, $I=L\sqcup M\sqcup R$. We will be looking for two disjoint sets $S_{1}\subseteq L\sqcup M$ and $S_{2}\subseteq M\sqcup R$ such that $|S_{1}\cap M|=|S_{2}\cap M|=|M|/4$ and $S_{1}\sqcup S_{2}$ is a solution. This problem can be solved by a reduction to the Weighted Orthogonal Vectors (WOV) problem that can be viewed as a hybrid of the Orthogonal Vectors and the $2$-SUM problems: given two families $\mathcal{L}$ and $\mathcal{R}$ of weighted sets and a target integer $t$, find two disjoint sets whose sum of weights is equal to $t$. A naive such reduction would proceed similarly to the algorithm by Horowitz and Sahni [15]: let
$\mathcal{L}$ =$=\{(S_{1}\cap M,w(S_{1}))\colon S_{1}\subseteq L\sqcup M,|S_{1}\cap M|=|M/4|\},$
$\mathcal{R}$ =$=\{(S_{2}\cap M,w(S_{2}))\colon S_{2}\subseteq M\sqcup R,|S_{2}\cap M|=|M/4|\}.$
Then, it remains to solve WOV for $\mathcal{L}$, $\mathcal{R}$ and $t$. The bad news is that $|L\sqcup M|>n/2$, hence enumerating all $S_{1}\subseteq L\sqcup M$ leads to a running time worse than $O^{*}(2^{n/2})$. The good news is that the solution $S$ is represented by exponentially many pairs $(S_{1},S_{2})$: there are many ways to distribute $S\cap M$ between $S_{1}$ and $S_{2}$. This allows one to use the representation technique: as in the algorithm by Howgrave-Graham and Joux [16], take a random prime $p$ of the order $2^{|M|/2}$ and a random remainder $r\in\mathbb{Z}_{p}$; filter $\mathcal{L}$ and $\mathcal{R}$ to leave only subsets $S_{1}$ and $S_{2}$ such that $w(S_{1})\equiv_{p}r$ and $w(S_{2})\equiv_{p}{t-r}$. This way, one reduces the space complexity since $\mathcal{L}$ and $\mathcal{R}$ become about $p$ times smaller.
To further reduce the space complexity of the algorithm and to simplify it at the same time, we use the following key idea. We use another prime $q$ to further filter $\mathcal{L}$ and $\mathcal{R}$: this allows us to reduce the original instance to a larger number of instances that in turn have smaller size. The idea is best illustrated by the following toy example. Consider an instance $I$ of size $n$ of Subset Sum. As in the algorithm by Horowitz and Sahni [15], partition $I$ arbitrarily into two parts $L$ and $R$ of size $n/2$. Assume now that we are given a magic integer $q=\Theta(2^{n/4})$ that hashes all subsets of $L$ and $R$ almost perfectly: for each $r\in\mathbb{Z}_{q}$,
|$|\{A\subseteq L\colon w(A)\equiv_{q}r\}|,\,|\{B\subseteq R\colon w(B)\equiv_{q}{t-r}\}|=O(2^{n/4})\,.$
This allows us to solve $I$ in time $O^{*}(2^{n/2})$ and space $O^{*}(2^{n/4})$ without using a reduction to $4$-SUM and priority queues machinery of Schroeppel and Shamir’s algorithm: for each $r$, construct sequences of all subsets of $L$ and $R$ having weights $r$ and $t-r$ modulo $q$, respectively, and solve $2$-SUM for two resulting sequences of length $O(2^{n/4})$. Thus, instead of reducing $I$ to a single instance of $2$-SUM of size $2^{n/2}$, we reduce it to $2^{n/4}$ instances of size $2^{n/4}$. We show that even though in general a randomly chosen integer $q$ does not give a uniform distribution of remainders, one can still use this idea to reduce the space.
On a technical side, we choose a mixer $M$ more carefully. This allows us to cover various corner cases, when the size of the solution $S$ is not $n/2$ and when the mixer $M$ is far from being perfect, in the same manner, thus simplifying the algorithm by Nederlof and Węgrzycki.
2 General Setting
2.1 Modular Arithmetic
We write $a\equiv_{m}b$ to denote that integers $a$ and $b$ have the same remainder modulo $m$. By $\mathbb{Z}_{m}$ we denote the ring of remainders modulo $m$. We make use of the following well known fact.
Theorem 2.1 (Chinese Remainder Theorem).
For any coprime positive integers $p$ and $q$ and any $a\in\mathbb{Z}_{p}$ and $b\in\mathbb{Z}_{q}$, there exists a unique $x\in\mathbb{Z}_{pq}$ such that $x\equiv_{p}a$ and $x\equiv_{q}b$. Furthermore, it can be computed (using the extended Euclidean algorithm) in time $O(\log^{2}p+\log^{2}q)$.
2.2 Prime Numbers
Given an integer $t$, one can generate a uniform random prime from $[t,2t]$ in time $O(\log^{O(1)}t)$: to do this, one selects a random integer from $[t,2t]$ and checks whether it is prime [18]; the expected number of probes is $O(\log t)$ due to the law of distribution of prime numbers. We also use the following estimate: any positive integer $k$ has at most $\log_{2}k$ distinct prime divisors; hence the probability that a random prime from $[t,2t]$ divides $k$ is at most $O(\frac{\log k\log t}{t})$.
2.3 Probability Amplification
We use the following standard success probability amplification trick: if a “good” event $A$ occurs with probability at least $p$, then for $A$ to occur at least once with constant probability, $1/p$ repetitions are sufficient: $(1-p)^{1/p}\leq(e^{-p})^{1/p}=e^{-1}\,.$
2.4 Growth Rate and Entropy
Much like $O(\cdot)$ hides constant factors, $O^{*}(\cdot)$ and $\widetilde{O}(\cdot)$ hide factors that grow polynomially in the input size and polylogarithmically, respectively: for example, $n^{2}\cdot 1.5^{n}=O^{*}(1.5^{n})$ and $\log^{3}n\cdot n^{2}=\widetilde{O}(n^{2})$. $\Omega^{*}(\cdot)$, $\Theta^{*}(\cdot)$, $\widetilde{\Omega}(\cdot)$, and $\widetilde{\Theta}(\cdot)$ are defined similarly.
We use the following well known estimates for binomial coefficients: for any constant $0\leq\alpha\leq 1$,
$\Omega(n^{-1/2})2^{h(\alpha)n}\leq\binom{n}{\alpha n}\leq 2^{h(\alpha)n}\text{and hence }\binom{n}{\alpha n}=\Theta^{*}\left(2^{h(\alpha)n}\right),$ (1)
where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ is the binary entropy function. The function $h$ is known to be concave: for any $0\leq\beta\leq 1$ and any $0\leq x,y\leq 1$, $\beta h(x)+(1-\beta)h(y)\leq h(\beta x+(1-\beta)y)\,.$
2.5 Sets and Sums
For a positive integer $n$, $[n]=\{1,2,\dotsc,n\}$. For disjoint sets $A$ and $B$, by $A\sqcup B$ we denote their union.
Let $S\subseteq\mathbb{Z}$ be a finite set of integers. By $w(S)$ we denote the weight of $S$: $w(S)=\sum_{a\in S}a$. By $2^{S}$ we denote the power set of $S$, i.e., the set of its subsets: $2^{S}=\{A\colon A\subseteq S\}$. By $\binom{S}{k}$ we denote the set of all subsets of $S$ of size $k$: $\binom{S}{k}=\{A\subseteq S\colon|A|=k\}$. For a family of sets $\mathcal{F}\subseteq 2^{S}$, by $w(\mathcal{F})$ we denote the set of their weights: $w(\mathcal{F})=\{w(A)\colon A\in\mathcal{F}\}$.
Given $S$, one can compute $|w(2^{S})|$ in time and space $O^{*}(2^{|S|})$ by iterating over all of its subsets. If $|w(2^{S})|=2^{(1-\varepsilon)|S|}$, we call $S$ an $\varepsilon$-mixer.
Lemma 2.1.
Let $S\subseteq A$ and $|A|=n$. For $m=O(1)$ uniformly randomly chosen disjoint non-empty sets $A_{1},\dotsc,A_{m}\subseteq A$,
$\Pr\left[\frac{|A_{i}\cap S|}{|A_{i}|}\approx\frac{|S|}{|A|}\text{, for all $i\in[m]$}\right]=\Omega^{*}(1)\;,$
where $\frac{|A_{i}\cap S|}{|A_{i}|}\approx\frac{|S|}{|A|}$ means that $\left||A_{i}\cap S|-|A_{i}|\frac{|S|}{|A|}\right|=O(1)$.
Proof.
We will show that
$\Pr\left[|A_{i}\cap S|=\left\lfloor|A_{i}|\frac{|S|}{|A|}\right\rfloor\,\forall i\in[m-1]\land|A_{m}\cap S|=\left\lfloor|A_{m}|\frac{|S|-\sum_{i<m}\lfloor A_{i}\frac{|S|}{|A|}\rfloor}{|A|-\sum_{i<m}|A_{i}|}\right\rfloor\right]=\Omega^{*}(1)\;.$
After that it is easy to see that $\left||A_{i}\cap S|-|A_{i}|\frac{|S|}{|A|}\right|=O(1),\;\forall i\in[m]$.
Hence, we can think that all numbers $|A_{i}|\frac{|S|}{|A|}$ are integers.
Since $m=O(1)$, it suffices to estimate the probability for $m=1$. Let $B\subseteq A$ be a random set and
$\gamma=\frac{\left\lfloor|B|\cdot\frac{|S|}{|A|}\right\rfloor}{|B|}.$
Then
$\Pr\left[|B\cap S|=\left\lfloor|B|\frac{|S|}{|A|}\right\rfloor\right]$ $=\frac{\binom{|S|}{\gamma|B|}\binom{|A\setminus S|}{(1-\gamma)|B|}}{\binom{|A|}{|B|}}$ (by (1))
$=\Omega^{*}\left(2^{\gamma|A|h(|B|/|A|)+(1-\gamma)|A|h(|B|/|A|)-|A|h(|B|/|A|)}\right)=\Omega^{*}(1).$
2.6 Weighted Orthogonal Vectors
Definition 2.1 (Weighted Orthogonal Vectors, WOV).
Given families of $N$ weighted sets $\mathcal{A},\mathcal{B}\subseteq 2^{[d]}\times\mathbb{N}$, and a target integer $t$, the problem is to find $(A,w_{A})\in\mathcal{A}$ and $(B,w_{B})\in\mathcal{B}$ such that $A\cap B=\emptyset$ and $w_{A}+w_{B}=t$.
Lemma 2.2.
For any $\sigma\in[0,\frac{1}{2}]$, there is a Monte-Carlo algorithm that, given $\mathcal{A}\subseteq\binom{[d]}{\sigma d}\times\mathbb{N}$, $\mathcal{B}\subseteq\binom{[d]}{(1/2-\sigma)d}{\times\mathbb{N}}$ and a target integer $t$, solves $\operatorname{WOV}(\mathcal{A},\mathcal{B})$ in time $\widetilde{O}\left((|\mathcal{A}|+|\mathcal{B}|)2^{d(1-h(1/4))}\right)$ and space $\widetilde{O}\left(|\mathcal{A}|+|\mathcal{B}|+2^{d}\right)$.
To prove the lemma, we utilize the algorithm from [23], but with a different parameter range.
Definition 2.2 (Orthogonal Vectors, OV).
Given families of vectors $\mathcal{A},\mathcal{B}\subseteq 2^{[d]}$. Decide if there exists a pair $A\in\mathcal{A},B\in\mathcal{B}$ such that $A\cap B=\emptyset$.
Lemma 2.3.
For any $\mu\in[0,\frac{1}{2}]$, there is a Monte-Carlo algorithm that, given $\mathcal{A}\subseteq\binom{d}{\mu d}$ and $\mathcal{B}\subseteq\binom{d}{(1/2-\mu)d}$, solves $\operatorname{OV}(\mathcal{A},\mathcal{B})$ in time
$\widetilde{O}\left((|\mathcal{A}|+|\mathcal{B}|)2^{d(1-h(1/4))}\right)$
and space $\widetilde{O}\left(|\mathcal{A}|+|\mathcal{B}|+2^{d}\right)$.
Proof.
In case $\mu\in[0.2,0.3]$, this becomes exactly Theorem 6.1 from [23] applied for $\alpha=\frac{1}{2},\sigma=2\mu$. The restriction $\sigma\in[0.4,0.6]$ in that theorem appears due to the following lemma.
Lemma 2.4 (Lemma B.1 from [23]).
For large enough $n$ and $\lambda\in[0.4,0.5]$ and $\sigma\in[0.4,0.6]$, the following inequality holds:
$\min_{x}\left\{\frac{\binom{1-\lambda\sigma}{x-\lambda\sigma}_{n}+\binom{1-(1-\sigma)\lambda}{x}_{n}}{\binom{1-\lambda}{x-\lambda\sigma}_{n}}\right\}\leq 2^{n(1/2+\lambda-h(\lambda/2))}n^{O(1)}.$
Here, $\binom{\alpha}{\beta}_{n}$ denotes $\binom{\alpha n}{\beta n}$.
We need that condition to hold for any $\sigma\in[0,1]$ when $\lambda=0.5$. For this reason, we adjust the statement of the lemma.
Lemma 2.5.
For large enough $n$ and $\sigma\in[0,1]$ the following inequality holds:
$\min_{x}\left\{\frac{\binom{1-\sigma/2}{x-\sigma/2}_{n}+\binom{1-(1-\sigma)/2}{x}_{n}}{\binom{1/2}{x-\sigma/2}_{n}}\right\}\leq 2^{n(1-h(1/4))}n^{O(1)}.$
Proof.
We will use $\mu=\frac{\sigma}{2}$, $\mu\in[0,0.5]$ to simplify the calculations.
$\min_{x}\left\{\frac{\binom{1-\mu}{x-\mu}_{n}+\binom{\frac{1}{2}+\mu}{x}_{n}}{\binom{1/2}{x-\mu}_{n}}\right\}=O^{*}(\min_{x}\{2^{\max\{(1-\mu)h(\frac{x-\mu}{1-\mu}),(\frac{1}{2}+\mu)h(\frac{2x}{1+2\mu})\}-\frac{1}{2}h(2x-2\mu)}\})\leq$
(we use $x=\frac{1}{2}+\log(3)(\mu-\frac{1}{4})$ as a good enough approximation of the minimum)
$\leq O^{*}(2^{max\{(1-\mu)h(\frac{(\frac{1}{2}+\log(3)(\mu-\frac{1}{4}))-\mu}{1-\mu}),(\frac{1}{2}+\mu)h(\frac{2(\frac{1}{2}+\log(3)(\mu-\frac{1}{4}))}{1+2\mu})\}-\frac{1}{2}h(2(\frac{1}{2}+\log(3)(\mu-\frac{1}{4}))-2\mu)})$
Let:
$f_{1}=(1-\mu)h\left(\frac{\left(\frac{1}{2}+\log(3)\left(\mu-\frac{1}{4}\right)\right)-\mu}{1-\mu}\right)-\frac{1}{2}h\left(2\left(\frac{1}{2}+\log(3)\left(\mu-\frac{1}{4}\right)\right)-2\mu\right)\;,$
$f_{2}=\left(\frac{1}{2}+\mu\right)h\left(\frac{2\left(\frac{1}{2}+\log(3)\left(\mu-\frac{1}{4}\right)\right)}{1+2\mu}\right)-\frac{1}{2}h\left(2\left(\frac{1}{2}+\log(3)\left(\mu-\frac{1}{4}\right)\right)-2\mu\right)\;.$
We will show that both of these functions have a maximum at $\mu=\frac{1}{4}$. Indeed,
$\frac{\partial^{2}f_{1}}{\partial\mu^{2}}=\frac{(3\log{3}-2)^{2}}{(1-\mu)(4\log{3}\mu-4\mu-\log{3}+2)(4\log(3)\mu-\log(3)-2)}\;,$
$\frac{\partial^{2}f_{2}}{\partial\mu^{2}}=-\frac{8(\log{3}-1)^{2}}{(4\log(3)\mu-4\mu-\log{3}+2)(4\log(3)\mu-4\mu-\log(3))}\;.$
By analyzing the roots of the polynomials in the denominators one can show that for $00\leq\mu\leq\frac{1}{2}:\frac{\partial^{2}f_{1}}{\partial\mu^{2}}<$ 0 and $\frac{\partial^{2}f_{2}}{\partial\mu^{2}}<0$. Which means that there is at most one maximum for each of the functions on this interval. Now we just need to notice that
$\frac{\partial f_{1}}{\partial\mu}|_{\mu=\frac{1}{2}}=\frac{\partial f_{2}}{\partial\mu}|_{\mu=\frac{1}{2}}=0$
and $f_{1}|_{mu=\frac{1}{2}}=f_{2}|_{mu=\frac{1}{2}}=1-h(\frac{1}{4})$. This concludes the proof. ∎
Now we can apply the proof of Theorem 6.1 from [23] with Lemma 2.4 replaced by Lemma 2.5. That completes the proof of Lemma 2.3. ∎
Proof of Lemma 2.2.
To prove the lemma we show a reduction from WOV to OV. We split $\mathcal{A}$ and $\mathcal{B}$ into the following groups of sets with the same sum in time $\widetilde{O}(|\mathcal{A}|+|\mathcal{B}|)$.
$\mathcal{L}(a)$ =$=\{A\colon(A,a)\in\mathcal{A}\}\;,$
$\mathcal{R}(b)$ =$=\{B\colon(B,b)\in\mathcal{B}\}\;.$
As in the algorithm for $2$-SUM, we use the two pointers technique to iterate over every possible $(a,b)$ such that $a+b=t$ in time $\widetilde{O}(\text{number of groups})$. For every such pair we call $\operatorname{OV}(\mathcal{L}(a),\mathcal{R}(b))$. By Lemma 2.3, its running time and space usage depend linearly on $(|\mathcal{L}(a)|+|\mathcal{R}(b)|)$, so the total time and space complexity are at most as for $\operatorname{OV}(\mathcal{A},\mathcal{B})$. ∎
3 Improved Space Upper Bound for Subset Sum
3.1 Algorithm
Let $(I,t)$ be an instance of Subset Sum (as usual, $n=|I|$) and assume that it is a yes-instance: we will present a randomized algorithm that is correct on no-instances with probability 1, so, as usual, the main challenge is to obtain $\Omega^{*}(1)$ success probability for yes-instances. Let $S\subseteq I$ be a solution: $w(S)=t$. It is unknown to us, but we may assume that we know its size: there are just $n$ possibilities for $|S|$, so with a polynomial overhead we can enumerate all of them. Thus, assume that $|S|=\alpha n$ where $0<\alpha\leq 1$ is known to us.
For a parameter $\beta=\beta(\alpha)<0.15$ to be specified later, select randomly pairwise disjoint sets $M_{L},M,M_{R}\in\binom{I}{\beta n}$. In time and space $O^{*}(2^{0.15n})$, find $\varepsilon_{L},\varepsilon,\varepsilon_{R}$ such that $M_{L}$ is an $\varepsilon_{L}$-mixer, $M$ is an $\varepsilon$-mixer, and $M_{R}$ is $\varepsilon_{R}$-mixer. Without loss of generality, we assume that $\varepsilon\leq\varepsilon_{L},\varepsilon_{R}$. Consider the probability that $S$ touches exactly half of $M_{L}\sqcup M\sqcup M_{R}$:
$\Pr\left[|(M_{L}\cup M\cup M_{R})\cap S|=3\beta n/2\right]=\frac{\binom{\alpha n}{3\beta n/2}\binom{(1-\alpha)n}{3\beta n/2}}{\binom{n}{3\beta n}}=\Theta^{*}\left(2^{-\lambda n}\right),$
where the value of $\lambda$ is estimated using (1):
$\frac{\binom{\alpha n}{3\beta n/2}\binom{(1-\alpha)n}{3\beta n/2}}{\binom{n}{3\beta n}}=2^{-\lambda n}$ $\iff\frac{1}{n}\log\frac{\binom{n}{3\beta n}}{\binom{\alpha n}{3\beta n/2}\binom{(1-\alpha)n}{3\beta n/2}}=\lambda\iff$
$\iff\lambda=h(3\beta)-\alpha h\left(\frac{3\beta}{2\alpha}\right)-(1-\alpha)h\left(\frac{3\beta}{2(1-\alpha)}\right).$ (2)
By repeating this process $2^{\lambda n}$ times, we ensure that the event $|(M_{L}\cup M\cup M_{R})\cap S|=3\beta n/2$ happens with probability $\Omega^{*}(1)$ (recall Section 2.3). Further, we may assume that
$|M_{L}\cap S|=|M\cap S|=|M_{R}\cap S|=\beta n/2,$
as by Lemma 2.1, $|M_{L}\cap S|=|M\cap S|=|M_{R}\cap S|=\beta n/2+O(1)$ with probability $\Omega^{*}(1)$ conditioned on the event that $S$ touches $23\beta n/$ 2 elements of $M_{L}\sqcup M\sqcup M_{R}$. An exact equality can be assumed since we can search for an exact value in the neighbourhood of $\beta n/2$ and the constant difference will not affect the memory and time complexity. In the following applications of Lemma 2.1, we omit $O(1)$ factor for similar considerations.
We argue that either $S\cap M$ or $M\setminus S$ is an $(\leq\varepsilon)$-mixer. Indeed, $M$ is an $\varepsilon$-mixer and hence $|w(2^{M})|=2^{(1-\varepsilon)|M|}$. Now, $|S\cap M|=|M\setminus S|=|M|/2$ and $|w(2^{S\cap M})|\cdot|w(2^{M\setminus S})|\geq|w(2^{M})|\,.$ Hence,
$\max\left\{|w(2^{S\cap M})|,|w(2^{M\setminus S})|\right\}\geq 2^{(1-\varepsilon)|M|/2}$
implying that either $S\cap M$ or $M\setminus S$ is indeed a $(\leq\varepsilon)$-mixer. In the following, we assume that it is $S\cap M$ that is $(\leq\varepsilon)$-mixer. To assume this without loss of generality, we run the final algorithm twice — for $(I,t)$ and $(I,w(I)-t)$.
For any (finite and non-empty) set $A$, there exists $1\leq k\leq|A|$, such that $|w(\binom{A}{k})|\geq|w(2^{A})|/|A|$ (by the pigeonhole principle). Since $|w(\binom{A}{k})|=|w(\binom{A}{|A|-k})|$, we can even assume that $21\leq k\leq|A|/$ 2. Consider the corresponding $k$ for the set $A=S\cap M$ and let $\mu=k/(2|S\cap M|)$ (hence, $0\leq\mu\leq{0.25})$. Then,
$\left|w\left(\binom{S\cap M}{2\mu|S\cap M|}\right)\right|\geq\frac{\left|w\left(2^{S\cap M}\right)\right|}{|S\cap M|}\,.$ (3)
Partition $I\setminus(M_{L}\sqcup M\sqcup M_{R})$ randomly into $L_{1}\sqcup L_{2}\sqcup L_{3}\sqcup L_{4}\sqcup R_{1}\sqcup R_{2}\sqcup R_{3}\sqcup R_{4}$ where the exact size of all eight parts will be specified later. Let $L=L_{1}\sqcup L_{2}\sqcup L_{3}\sqcup L_{4}$ and $R=R_{1}\sqcup R_{2}\sqcup R_{3}\sqcup R_{4}$. By Lemma 2.1, with probability $\Omega^{*}(1)$ the set $S$ covers the same fraction of each of $L_{i}$’s and $R_{i}$’s. We denote this fraction by $\gamma$:
$\gamma=\frac{|(L\sqcup R)\cap S|}{|L\sqcup R|}=\frac{\alpha-3\beta/2}{1-3\beta}\,.$ (4)
Now, we apply the representation technique. Note that there exist many ways to partition $S$ into $S_{L}\sqcup S_{R}$ such that
$S_{L}\subseteq M_{L}\sqcup L\sqcup M\text{ and }S_{R}\subseteq M\sqcup R\sqcup M_{R}.$ (5)
Indeed, the elements from $S\cap M$ can be distributed arbitrary between $S_{L}$ and $S_{R}$. Below, we show that for a large random prime number $p$ and a random remainder $r$ modulo $p$, the probability that there exists a partition $S=S_{L}\sqcup S_{R}$ such that $w(S_{L})\equiv_{p}r$ is $\Omega^{*}(1)$ (informally, at least one representative $S_{L}\sqcup S_{R}$ of the original solution $S$ survives, even if we “hit” all such partitions with a large prime).
Lemma 3.1.
With constant probability over choices of a prime number $p$ such that $p\in[X/2,X]$, where $X\leq|w(2^{S\cap M})|$, the number of remainders $a\in\mathbb{Z}_{p}$ such that there exists a subset $M^{\prime}\subseteq S\cap M$ with $|M^{\prime}|=2\mu|S\cap M|$ and $w(M^{\prime})\equiv_{p}a$ is close to $p$:
$\Pr\left[\left|\left\{a\in\mathbb{Z}_{p}\colon\text{there exists }M^{\prime}\subseteq S\cap M,|M^{\prime}|=2\mu|S\cap M|,w(M^{\prime})\equiv_{p}a\right\}\right|\geq\Omega\left(p/n^{2}\right)\right]\geq 9/10.$
Proof of Lemma 3.1.
Let $Q=M\cap S$. Then $|Q|=\frac{\beta}{2}n$. Recall that $|w(\binom{Q}{2\mu|Q|})|\geq|w(2^{Q})|/|Q|$ and $p$ is a random prime in $[X/2,X]$.
Let $\mathcal{M}=\binom{Q}{2\mu|Q|}$, $c_{i}=|\{M\colon M\in\mathcal{M},w(M)\equiv_{p}i\}|$, and $d_{i}=[c_{i}>0]$. We need to show that
$\Pr[|\{i\colon d_{i}=1\}|\geq\Omega(p/n^{2})]\geq 9/10.$
The expected number of collisions is
$\mathbb{E}\left[\sum_{i}c_{i}^{2}\right]=\sum_{M_{1},M_{2}\in\mathcal{M}}\Pr\left[p\text{ divides }w(M_{2})-w(M_{1})\right]\leq|\mathcal{M}|+O(n^{2}|\mathcal{M}|^{2}/X).$
The last inequality follows from Section 2.2, all numbers are not greater than $2^{O(n)}$. Hence,
$\sum_{i}c_{i}^{2}=O(|\mathcal{M}|+n^{2}|\mathcal{M}|^{2}/X)=O(n^{2}|\mathcal{M}|^{2}/p)$
with constant probability. Then, the Cauchy–Schwartz inequality implies that
$|\mathcal{M}|=\sum_{i}c_{i}=\sum_{i}c_{i}d_{i}\leq\sqrt{\sum_{i}d_{i}^{2}}\sqrt{\sum_{i}c_{i}^{2}}=\sqrt{\sum_{i}d_{i}}\sqrt{\sum_{i}c_{i}^{2}}\,.$
Hence,
$\sum_{i}d_{i}\geq\frac{(\sum_{i}c_{i})^{2}}{\sum_{i}c_{i}^{2}}\geq\frac{|\mathcal{M}|^{2}}{O(n^{2}|\mathcal{M}|^{2}/p)}\geq\Omega(p/n^{2})\,.$
Take a random prime $2^{\frac{1-\varepsilon}{2}\beta n-1}\leq p\leq 2^{\frac{1-\varepsilon}{2}\beta n}$ and a random remainder $r\in\mathbb{Z}_{p}$. Lemma 3.1 ensures that with probability $\Omega^{*}(1)$, there exists $M^{\prime}\subseteq S\cap M$ such that $|M^{\prime}|=2\mu|S\cap M|=\mu\beta n$ and $w(M^{\prime})\equiv_{p}r$.
To find the partition $S_{L}\sqcup S_{R}$, we are going to enumerate all sets $A\subseteq M_{L}\sqcup L\sqcup M$ with $w(A)\equiv_{p}r$ and $|A\cap(M\cap S)|=\mu\beta n$ as well as all sets $B\subseteq M\sqcup R\sqcup M_{R}$ with $w(B)\equiv_{p}t-r$ and $|B\cap(M\cap S)|=(1/2-\mu)\beta n$. Then, it suffices to solve WOV for the sets of $A$’s and $B$’s.
To actually reduce to WOV, we first construct set families $\mathcal{M}_{L}\subseteq 2^{M_{L}}$ and $\mathcal{M}_{R}\subseteq 2^{M_{R}}$ such that $|\mathcal{M}_{L}|=|w(2^{M_{L}})|$ and $|\mathcal{M}_{R}|=|w(2^{M_{R}})|$. That is, $\mathcal{M}_{L}$ contains a single subset of $M_{L}$ for each possible weight. One can construct $\mathcal{M}_{L}$ (as well as $\mathcal{M}_{R}$) simply by going through all subsets of $M_{L}$ and checking, for each subset, whether its sum is already in $\mathcal{M}_{L}$.
Then, construct the following families (see Figure 1).
$\mathcal{Q}_{1}$ =$=\{M^{\prime}\cup L^{\prime}\colon M^{\prime}\in\mathcal{M}_{L},L^{\prime}\subseteq L_{1},|L^{\prime}|=\gamma|L_{1}|\},$ (6)
$\mathcal{Q}_{2}$ =$=\{L^{\prime}\colon L^{\prime}\subseteq L_{2},|L^{\prime}|=\gamma|L_{2}|\},$ (7)
$\mathcal{Q}_{3}$ =$=\{L^{\prime}\colon L^{\prime}\subseteq L_{3},|L^{\prime}|=\gamma|L_{3}|\},$ (8)
$\mathcal{Q}_{4}$ =$=\{L^{\prime}\cup M^{\prime}\colon L^{\prime}\subseteq L_{4},M^{\prime}\subseteq M,|L^{\prime}|=\gamma|L_{4}|,|M^{\prime}|=\mu\beta n\},$ (9)
$\mathcal{Q}_{1}^{\prime}$ =$=\{M^{\prime}\cup R^{\prime}\colon M^{\prime}\in\mathcal{M}_{R},R^{\prime}\subseteq R_{1},|R^{\prime}|=\gamma|R_{1}|\},$ (10)
$\mathcal{Q}_{2}^{\prime}$ =$=\{R^{\prime}\colon R^{\prime}\subseteq R_{2},|R^{\prime}|=\gamma|R_{2}|\},$ (11)
$\mathcal{Q}_{3}^{\prime}$ =$=\{R^{\prime}\colon R^{\prime}\subseteq R_{3},|R^{\prime}|=\gamma|R_{3}|\},$ (12)
$\mathcal{Q}_{4}^{\prime}$ =$=\{R^{\prime}\cup M^{\prime}\colon R^{\prime}\subseteq R_{4},M^{\prime}\subseteq M,|R^{\prime}|=\gamma|R_{4}|,|M^{\prime}|=(1/2-\mu)\beta n\}.$ (13)
[…]
Figure 1: Partition of the instance $I$ (white) and its solution $S$ (gray) into parts.
Let $\mathcal{Q}_{max}$ stand for $\mathcal{Q}_{i}$ or $\mathcal{Q}_{i}^{\prime}$ with the maximum size. For $a,p\in\mathbb{Z}_{>0}$, let
$\mathcal{X}_{a}=\left\{(Q_{1},Q_{2},Q_{3},Q_{4})\colon Q_{i}\in\mathcal{Q}_{i},\sum w(Q_{i})=a\right\}\text{ and }\mathcal{X}_{a,p}=\bigcup\limits_{a^{\prime}\equiv_{p}a}\mathcal{X}_{a^{\prime}}.$
Then, it remains to solve WOV for $\mathcal{X}_{r,p}$ and $\mathcal{X^{\prime}}_{(t-r),p}$. To do this, we use the idea used in the Arthur–Merlin protocol: we reduce to 4-SUM and use a cutoff for the number of found candidates. The following lemmas estimate the size of these set families.
Lemma 3.2.
With probability $\Omega^{*}(1)$ over the choice of $r$, the size of $\mathcal{X}_{r,p}$ is at most $O^{*}(|\mathcal{Q}_{max}|^{4}/p)$ and there exists $M_{1}\subseteq M\cap S$ such that $w((S\cap(L\cup M_{L}))\cup M_{1})\equiv_{p}r$.
Proof.
Let $A$ be an event of existing $M_{1}\subseteq M\cap S$ such that $w((S\cap(L\cup M_{L}))\cup M_{1})\equiv_{p}r$. Note that
$\mathbb{E}[\mathcal{X}_{r,p}]\leq|\mathcal{Q}_{max}|^{4}/p\text{ and }\Pr[\mathcal{X}_{r,p}>x]\leq\frac{|\mathcal{Q}_{max}|^{4}}{px}\,.$
As $\Pr[A]=\Omega^{*}(1)$, we can take $x=O^{*}(\frac{|\mathcal{Q}_{max}|^{4}}{p})$ in a way that will make $\Pr[\mathcal{X}_{r,p}>x]\leq\frac{\Pr[A]}{2}$. Then
$\Pr[\mathcal{X}_{r,p}\leq x\wedge A]\geq\Pr[A]-\Pr[\mathcal{X}_{r,p}>x]\geq\frac{\Pr[A]}{2}=\Omega^{*}(1)\,.$
To further reduce the size of these families, we choose another random prime number
$q=\Theta^{*}\left(\frac{2^{n/2}}{2^{\lambda n}|\mathcal{Q}_{max}|^{2}}\right)$
and iterate over all possible remainders $s\in\mathbb{Z}_{q}$. We aim to find two families of sets as the above ones, but with weights congruent to $w(S_{L})$ and $(t-w(S_{L}))$, respectively, modulo $p$ and $q$. While iterating on $s$, at some point we try $s\equiv_{q}w(S_{L})$, so we can assume that we guessed $s$ correctly and construct the families for the fixed $r$ and $s$. We now focus on the first family only, since we can then find the second one independently using the same algorithm.
For $a\in\mathbb{Z}_{>0}$, let $\mathcal{L}_{a}=\{(a,Q_{4}\cap M)\colon(Q_{1},Q_{2},Q_{3},Q_{4})\in\mathcal{X}_{a}\}$. For $a_{1},a_{2},p,q\in\mathbb{Z}_{>0}$, let
$\mathcal{X}_{a_{1},p,a_{2},q}=\bigcup\limits_{\begin{subarray}{c}a\equiv_{p}a_{1}\\ a\equiv_{q}a_{2}\end{subarray}}\mathcal{X}_{a}\text{ and }\mathcal{L}_{a_{1},p,a_{2},q}=\bigcup\limits_{\begin{subarray}{c}a\equiv_{p}a_{1}\\ a\equiv_{q}a_{2}\end{subarray}}\mathcal{L}_{a}.$
Our goal is to find a family $\mathcal{L}$ containing $\mathcal{L}_{w(S_{L})}$ in order to guarantee that we later find the solution using the WOV algorithm. But we need $\mathcal{L}$ to be small enough. Let $\ell=\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\max\{\mu,1/2-\mu\}|M|}$ represent an upper bound on the sum of the average size of $\mathcal{X}_{a_{1},p,a_{2},q}$ and the size of $\mathcal{L}_{w(S_{L})}$. We want $\mathcal{L}$ to be of size $\widetilde{O}(\ell)$. The following lemma shows that $\mathcal{X}_{w(S_{L}),p,w(S_{L}),q}$ contains not too many redundant sets.
Lemma 3.3.
With probability $\Omega^{*}(1)$ over the choice of $q$, $|\mathcal{X}_{r,p,w(S_{L}),q}\setminus\mathcal{X}_{w(S_{L})}|\leq\ell.$
Proof.
We estimate the expected number of elements in $\mathcal{\mathcal{missing}}{X}_{r,p,w(S_{L}),q}\setminus\mathcal{X}_{w(S_{L})}$. To do that, we calculate the sum of $\Pr_{q}[\sum w(Q_{i})\equiv_{q}w(S_{L})]$ over all tuples $(Q_{1},Q_{2},Q_{3},Q_{4})\in\mathcal{X}_{r,p}\setminus\mathcal{X}_{w(S_{L})}$ such that $\sum w(Q_{i})\neq w(S_{L})$.
Since $\Pr_{q}[\sum w(Q_{i})\equiv_{q}w(S_{L})]=O^{*}(\frac{1}{q})$ (recall Section 2.2),
$\mathbb{E}|\mathcal{\mathcal{missing}}{X}_{r,p,w(S_{L}),q}\setminus\mathcal{X}_{w(S_{L})}|=|\mathcal{X}_{r,p}\setminus\mathcal{X}_{w(S_{L})}|\cdot O^{*}(1).$
By Lemma 3.2, it is equal to $\widetilde{O}({|\mathcal{Q}_{max}|^{4}}/(pq))=\widetilde{O}(\ell)$. ∎
Let us find the first $\leq 2\ell+1$ elements of $\mathcal{X}_{r,s}$ using the algorithm from Lemma 3.4 which runs in time $\widetilde{O}(|\mathcal{Q}_{max}|^{2}+\ell)$ and uses $\widetilde{O}(|\mathcal{Q}_{max}|+\ell)$ space.
Lemma 3.4.
There is an algorithm that given $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$, $(r,p,s,q)$ and integer $m$, finds $\min(|\mathcal{X}_{r,p,s,q}|,m)$ elements of $\mathcal{X}_{r,p,s,q}$ using $\widetilde{O}(|\mathcal{Q}_{max}|^{2}+m)$ time and $\widetilde{O}(|\mathcal{Q}_{max}|+m)$ space.
Proof.
Let $v\in[0,pq-1]$ be an integer such that $v\equiv_{p}s$ and $v\equiv_{q}s$. We obtain such $v$ in $\Omega^{*}(1)$ time by Theorem 2.1. Our goal is to output every tuple $(Q_{1},Q_{2},Q_{3},Q_{4})$ such that $Q_{i}\in\mathcal{Q}_{i}$ and $\sum w(Q_{i})\equiv_{pq}v$, one by one and halt the algorithm when the size of the output becomes $m$.
We can assume that all weights of sets from $\mathcal{Q}_{i}$ are in $[0,pq-1]$. Then, we want to output the tuples with total weights equal to $v$, $v+pq$, $v+2pq$ or $v+3pq$. Now we run Schroeppel and Shamir’s [25] algorithm four times keeping the total size of the output at most $m$. ∎
Assume that $s\equiv_{q}w(S_{L})$. If the algorithm yields at most $2\ell$ elements then we know the whole $\mathcal{X}_{r,p,w(S_{L}),q}$, so we can now just construct $\mathcal{L}_{w(S_{L})}$ in $\widetilde{O}(\ell)$ time and space. Otherwise, by Lemma 3.3, we have that $|\mathcal{X}_{w(S_{L})}|>\ell$ with probability $\Omega^{*}(1)$. Since strictly more than half of the outputted tuples have weight exactly $w(S_{L})$, we can determine $w(S_{L})$ in $\widetilde{O}(\ell)$ time. When we know $w(S_{L})$, we use Lemma 3.5 to obtain $\mathcal{L}_{w(S_{L})}$ using $\widetilde{O}(|\mathcal{Q}_{max}|^{2})$ time and $\widetilde{O}(|\mathcal{Q}_{max}|)$ space.
Lemma 3.5.
There is an algorithm that given $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$ and $a$, finds $\mathcal{L}_{a}$ using $\widetilde{O}(|\mathcal{Q}_{max}|^{2})$ time and $\widetilde{O}(|\mathcal{Q}_{max}|)$ space.
Proof.
Schroeppel and Shamir [25] use priority queues to implement a data structure $D_{1}$ for the following task: output all elements from $\{w(Q_{1})+w(Q_{2})\colon Q_{1}\in\mathcal{Q}_{1},Q_{2}\in\mathcal{Q}_{2}\}$ in non-decreasing order using $\widetilde{O}(|\mathcal{Q}_{1}|\cdot|\mathcal{Q}_{2}|)$ time and $\widetilde{O}(|\mathcal{Q}_{1}|+|\mathcal{Q}_{2}|)$ space. Similarly, a data structure $D_{2}$ outputs all elements from $\{(w(Q_{3})+w(Q_{4}),Q_{4}\cap M)\colon Q_{3}\in\mathcal{Q}_{3},Q_{4}\in\mathcal{Q}_{4}\}$ in non-increasing order (we compare elements by their first coordinate) using $\widetilde{O}(|\mathcal{Q}_{3}|\cdot|\mathcal{Q}_{4}|)$ time and $\widetilde{O}(|\mathcal{Q}_{3}|+|\mathcal{Q}_{4}|)$ space. By $\operatorname{inc}(w(\mathcal{Q}_{1},\mathcal{Q}_{2}))$ and $\operatorname{dec}(w(\mathcal{Q}_{3},\mathcal{Q}_{4}))$ we denote initilization methods for these data structures, by $\operatorname{pop}()$ we denote their method that gives the next element.
Algorithm 1 solves the problem using data structures $D_{1}$ and $D_{2}$. It works in time $\widetilde{O}(|\mathcal{Q}_{1}|\cdot|\mathcal{Q}_{2}|+|\mathcal{Q}_{3}|\cdot|\mathcal{Q}_{4}|)$, uses $\widetilde{O}(|\mathcal{Q}_{1}|+|\mathcal{Q}_{2}|+|\mathcal{Q}_{3}|+|\mathcal{Q}_{4}|)$ space, and outputs $|\mathcal{L}_{a}|$ elements. Since the number of outputted subsets is at most $|\mathcal{Q}_{4}|$, even if we store the output, we still need only $\widetilde{O}(|\mathcal{Q}_{max}|^{2})$ time and $\widetilde{O}(|\mathcal{Q}_{max}|)$ space.
Input : $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4},a$
Output : $\mathcal{L}_{a}=\{(a,Q_{4}\cap M)\colon(Q_{1},Q_{2},Q_{3},Q_{4})\in\mathcal{X}_{a}\}$
1 Initialize $D_{1}$ = inc($w(\mathcal{Q}_{1},\mathcal{Q}_{2})$)
2 Initialize $D_{2}$ = dec($w(\mathcal{Q}_{3},\mathcal{Q}_{4})$)
3 while $(d_{2},A)=D_{2}$.pop() do
4       while $d_{1}=D_{1}$.pop() and $d_{1}+d_{2}<a$ do
5             skip $d_{1}$
6      if $d_{1}+d_{2}=a$ and not used[$A$] then
7             output($(a,A)$)
8             used[$A$] = True
Algorithm 1 Pseudocode for Lemma 3.5.
Note that the size of $\mathcal{L}_{a}$ is at most $\binom{|M|}{\mu|M|}$. Summarizing, we get the following.
Lemma 3.6.
There is an algorithm that given $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$ and $(r,p,s,q)$, finds $\mathcal{L}$ using $\widetilde{O}\left(|\mathcal{Q}_{max}|^{2}+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\mu|M|}\right)$ time and $\widetilde{O}\left(|\mathcal{Q}_{max}|+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\mu|M|}\right)$ space. And $|\mathcal{L}|=\widetilde{O}\left(\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\mu|M|}\right)$.
Similarly, using Lemma 3.6, we find subset $\mathcal{R}$ for the right side. We call $\mathcal{L},\mathcal{R}$) and return True if it finds a solution. The WOV algorithm works in time $\widetilde{O}((|\mathcal{L}|+|\mathcal{R}|)2^{\beta n(1-h(1/4)}))$ and space $\widetilde{O}(|\mathcal{L}|+|\mathcal{R}|+2^{\beta n})$ by Lemma 2.2.
It is worth noting that for most choices of $s$ our assumption that $s\equiv_{q}w(S_{L})$ does not hold. Also, with probability $1-\Omega^{*}(1)$, $|\mathcal{X}_{r,p,w(S_{L}),q}\setminus\mathcal{X}_{w(S_{L})}|$ may happen to be greater than $\ell$. If any of these cases occurs, we may find a set $\mathcal{L}$ that was not intended, or we may even not find anything at all (if after applying the algorithm from Lemma 3.4 we obtain $12\ell+$ 1 elements, each of which occurs less than $\ell+1$ times). In the second scenario, we simply continue with the next choice of $s$. Otherwise, we cannot determine whether we have found the correct $\mathcal{L}$ or not. However, note that Lemma 3.6 holds anyway. Also, if an incorrect $\mathcal{L}$ is found, the outcome of our algorithm remains unaffected. If $\mathcal{L},\mathcal{R}$) finds a solution, it is a solution for the original problem; if not, we simply proceed to the next iteration of the loop in line 2.
Now we are ready to present the final Algorithm 2. We will choose the exact values for its parameters later.
Input : $(I,t)$.
Output : True, if $S\subseteq I$ with $w(S)=t$ exists, otherwise False.
1
2 repeat $2^{\lambda n}$ times
3       Select random disjoint subsets $M_{L},M,M_{R}\subseteq I$ of size $\beta n$
4       foreach $\mu$ such that $\mu\beta n\in[0,\beta n/2]$ do
5             Randomly partition $I\setminus(M_{L}\sqcup M\sqcup M_{R})$ into $L,R$ with size as in (15)–(16)
6             Pick a prime $p=\Theta(2^{(1-\varepsilon)\beta n/2})$ and $r\in\mathbb{Z}_{p}$ at random
7             Construct $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$ and $\mathcal{Q}_{1}^{\prime},\mathcal{Q}_{2}^{\prime},\mathcal{Q}_{3}^{\prime},\mathcal{Q}_{4}^{\prime}$ as defined in (6)–(13)
8             Pick a random prime $q=\Theta^{*}\left({2^{n/2}}/(2^{\lambda n}|\mathcal{Q}_{max}|^{2})\right)$
9
10            foreach $s\in\mathbb{Z}_{q}$ do
11                   Construct $\mathcal{L}$ and $\mathcal{R}$ (as in Lemma 3.6)
12                   if $\mathcal{L},\mathcal{R}$) (as in Lemma 2.2) then
13                         return True
return False
Algorithm 2 The main algorithm.
3.2 Correctness
Lemma 3.7.
The success probability of the algorithm is $\Omega^{*}(1)$.
Proof.
To detect that $I$ is indeed a yes-instance, an algorithm iteration relies on a number of events. The first one is $|(M_{L}\cup M\cup M_{R})\cap S|=3\beta n/2$ ,that occurs with probability $\Theta^{*}\left(2^{-\lambda n}\right)$. All the others occur with probability $\Omega^{*}(1)$, conditioned on the event that all the previous events occur. This implies that all of them occur with probability $\Omega^{*}\left(2^{-\lambda n}\right)$, and after repeating it $2^{\lambda n}$ times, the success probability of the resulting algorithm is $\Omega^{*}(1)$. Below, we recall all the considered events and provide links to lemmas proving their conditional probabilities.
  • $|(M_{L}\cup M\cup M_{R})\cap S|=3\beta n/2$, see (2).
  • $|M_{L}\cap S|=|M\cap S|=|M_{R}\cap S|=\frac{\beta}{2}n$, see Lemma 2.1.
  • $|S\cap L_{i}|=\gamma|L_{i}|$ and $|S\cap R_{i}|=\gamma|R_{i}|$, see Lemma 2.1.
  • $p$ is a prime number, see Section 2.2
  • |$\left|\left\{a\in\mathbb{Z}_{p}\colon\exists M^{\prime}\subseteq S\cap M,|M^{\prime}|=2\mu|S\cap M|,w(M^{\prime})\equiv_{p}a\right\}\right|\geq\Omega\left(p/n^{2}\right)$, see Lemma 3.1.
  • $|\mathcal{X}_{r,p}|=O^{*}(|\mathcal{Q}_{max}|^{4}/p)$ and there exists a partition $M_{1}\sqcup M_{2}=S\cap M$ such that $w(S\cap(M_{L}\cup L\cup M_{1}))\equiv_{p}r$ and $w(S\cap(M_{2}\cup R\cup M_{R}))\equiv_{p}t-r$, see Lemma 3.2.
  • $q$ is a prime number, see Section 2.2
  • $|\mathcal{X}_{r,p,w(S_{L}),q}\setminus\mathcal{X}_{w(S_{L})}|\leq\ell$, see Lemma 3.3.
In case any of those events does not happen, the algorithm completes the iteration without affecting the final result. ∎
3.3 Setting the Parameters
Recall that $\alpha=\frac{|S|}{n}$ and we iterate over all possible $\alpha$. We define $\beta=\beta(\alpha)$ as follows:
$\beta(\alpha)=\begin{cases}0.13,&\alpha\in[0.45,0.55]\\ 0,&\text{otherwise}\end{cases}$ (14)
Let us consider an iteration of the loop on line 2. To this point, we know the values of $\varepsilon$, $\varepsilon_{L}$, $\varepsilon_{R}$ and $\mu$, so we consider them as fixed numbers. Since $|M_{L}|=|M|=|M_{R}|=\beta n$, we have that $|I\setminus(M_{L}\sqcup M\sqcup M_{R})|=(1-3\beta)n$. Let us consider a partition $I\setminus(M_{L}\sqcup M\sqcup M_{R})=L_{1}\sqcup L_{2}\sqcup L_{3}\sqcup L_{4}\sqcup R_{1}\sqcup R_{2}\sqcup R_{3}\sqcup R_{4}$ and estimate the size of $\mathcal{Q}_{i}$ and $\mathcal{Q}^{\prime}_{i}$. Recall that $\gamma=\frac{|(L\cup R)\cap S|}{|L\cup R|}=\frac{\alpha-\frac{3}{2}\beta}{1-3\beta}$. We use $\varepsilon$ instead of $\varepsilon_{L}$ and $\varepsilon_{R}$ since $\varepsilon\leq\varepsilon_{L},\varepsilon_{R}$.
$|\mathcal{Q}_{1}|$ $\leq\binom{|L_{1}|}{\gamma|L_{1}|}\cdot 2^{(1-\varepsilon)\beta n}$ $|\mathcal{Q}_{1}^{\prime}|$ $\leq\binom{|R_{1}|}{\gamma|R_{1}|}\cdot 2^{(1-\varepsilon)\beta n}$
$|\mathcal{Q}_{2}|$ $=\binom{|L_{2}|}{\gamma|L_{2}|}$ $|\mathcal{Q}_{2}^{\prime}|$ $=\binom{|R_{2}|}{\gamma|R_{2}|}$
$|\mathcal{Q}_{3}|$ $=\binom{|L_{3}|}{\gamma|L_{3}|}$ $|\mathcal{Q}_{3}^{\prime}|$ $=\binom{|R_{3}|}{\gamma|R_{3}|}$
$|\mathcal{Q}_{4}|$ $=\binom{|L_{4}|}{\gamma|L_{4}|}\cdot\binom{|M|}{\mu\beta n}$ $|\mathcal{Q}_{4}^{\prime}|$ $=\binom{|R_{4}|}{\gamma|R_{4}|}\cdot\binom{|M|}{\left(\frac{1}{2}-\mu\right)\beta n}$
We construct $\mathcal{L}$ and $\mathcal{R}$ in time depending on $|\mathcal{Q}_{max}|^{2}+|\mathcal{Q}^{\prime}_{max}|^{2}$. Note that this function is minimized when all $|\mathcal{Q}_{i}|$ and $|\mathcal{Q}^{\prime}_{i}|$ are roughly the same. We can achieve that by a careful choice of $|L_{i}|$ and $|R_{i}|$.
Let us define sizes of $|L_{i}|$ and $|R_{i}|$ and show that all the parameters are chosen correctly, so we never, for example, choose the size for $L_{i}$ to be negative, or repeat the algorithm a number of times smaller than one.
Lemma 3.8.
$\lambda$ defined in Equation (2) is a non-negative number.
Proof.
Entropy is a concave function, hence
$\alpha h\left(\frac{3\beta}{2\alpha}\right)+(1-\alpha)h\left(\frac{3\beta}{2(1-\alpha)}\right)\leq h\left(\frac{3\beta}{2}+\frac{3\beta}{2}\right)=h(3\beta)\;.$
This implies that $\lambda\geq 0$. ∎
Now we set $|L|$ and $|R|$ as the following.
$|L|$ $=\frac{1-3\beta-\chi\beta}{2}n$ (15)
$|R|$ $=\frac{1-3\beta+\chi\beta}{2}n$ (16)
where $\chi=\frac{h(\mu)-h\left(\frac{1}{2}-\mu\right)}{h(\gamma)}$ is a balancing parameter needed to ensure that the proportion $\mu$ of elements of $M\cap S$ that go to the left side is not equal to $|M\cap S|/2$.
Lemma 3.9.
For our choice of $\beta$, for every $\alpha\in[0,1]$, $|L|\geq 0$, where
$|L|=n\frac{h(\gamma)-3\beta h(\gamma)-\beta h(\mu)+\beta h(1/2-\mu)}{2h(\gamma)}\;.$
Proof.
$|L|\geq 0\iff h(\gamma)-3\beta h(\gamma)+\beta(-h(\mu)+h(1/2-\mu))\geq 0\\ h(\gamma)-3\beta h(\gamma)+\beta(\underbrace{-h(\mu)+h(1/2-\mu)}_{\geq-1})\geq h(\gamma)-3\beta h(\gamma)-\beta=h(\gamma)(1-3\beta)-\beta$
Consider two cases:
  1. 1.
    $\alpha\in[0.45,0.55]$, then $\beta=0,h(\gamma)\geq h(0.45)\geq 0.99$ and
    $h(\gamma)(1-3\beta)-\beta\geq 0.99(1-3\cdot 0.13)-0.13\geq 0.47$
  2. 2.
    $\alpha\in[0,0.45)\cup(0.55,1]$, then $\beta=0$ and $|L|=\frac{1}{2}n$.
Lemma 3.10.
For our choice of $\beta$, for every $\alpha\in[0,1]$ we can choose $|L_{1}|$, $|L_{2}|$, $|L_{3}|$, $|L_{4}|$, $|R_{1}|$, $|R_{2}|$, $|R_{3}|$, $|R_{4}|$ $\geq 0$ such that $|L_{1}|+|L_{2}|+|L_{3}|+|L_{4}|=|L|$, $|R_{1}|+|R_{2}|+|R_{3}|+|R_{4}|=|R|$, and
$\log|\mathcal{Q}_{max}|$ $\leq n\cdot\max\{\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)+\beta h\left(1/2-\mu\right)}{8},\;(1-\varepsilon)\beta\}\;.$ (17)
Proof.
To make all $|\mathcal{Q}_{i}|$ and $|\mathcal{Q}_{i}^{\prime}|$ equal $n\frac{(h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)+\beta h\left(1/2-\mu\right))}{8}$ we set sizes of $L_{i}$ and $R_{i}$ as follows:
$|L_{1}|$ $=n\frac{h(\gamma)-3\beta h(\gamma)-6\beta+6\varepsilon\beta+\beta h(\mu)+\beta h\left(\frac{1}{2}-\mu\right)}{8h(\gamma)}=|R_{1}|$ (18)
$|L_{2}|=|L_{3}|$ $=n\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h\left(\mu\right)+\beta h\left(\frac{1}{2}-\mu\right)}{8h(\gamma)}=|R_{2}|=|R_{3}|$ (19)
$|L_{4}|$ $=n\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta-7\beta h(\mu)+\beta h\left(\frac{1}{2}-\mu\right)}{8h(\gamma)}$ (20)
$|R_{4}|$ $=n\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)-7\beta h\left(\frac{1}{2}-\mu\right)}{8h(\gamma)}$ (21)
If all $|L_{i}|$ and $|R_{i}|$ are non-negative, then the statement of the lemma holds. Assume that some values are negative. In this case we replace all $|L_{i}|<0$ with $|L_{i}|=0$ and scale all the $|L_{i}|$ in such a way that $|L_{1}|+|L_{2}|+|L_{3}|+|L_{4}|=|L|$. Then we do the same procedure with $|R_{i}|$. We can see that for every $|L_{i}|=0$, $|\mathcal{Q}_{i}|\leq 2^{(1-\varepsilon)\beta n}$, and for $|L_{i}|>0$, $|\mathcal{Q}_{i}|\leq n\frac{(h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)+\beta h\left(1/2-\mu\right))}{8}$. The same works for $|\mathcal{Q}^{\prime}_{i}|$. Thus, the lemma holds. ∎
We have set all the parameters and now we are ready to estimate the time and space complexity of Algorithm 2.
3.4 Time Complexity
The running time is dominated by the following parts.
  • Repeating the whole algorithm $2^{\lambda n}$ times (Line 2).
  • Constructing $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$ and $\mathcal{Q}_{1}^{\prime},\mathcal{Q}_{2}^{\prime},\mathcal{Q}_{3}^{\prime},\mathcal{Q}_{4}^{\prime}$ takes $O^{*}(|\mathcal{Q}_{max}|+2^{\beta n})$ time (Line 2).
  • Repeating $q$ times (Line 2).
  • Finding $\mathcal{L},\mathcal{R}$ using Lemma 3.6 in time $O^{*}\left(|\mathcal{Q}_{max}|^{2}+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\max\{\mu,1/2-\mu\}|M|}\right)$ (Line 2).
  • Solving $(\mathcal{L},\mathcal{R})$ using Lemma 2.2 in time $O^{*}\left((|\mathcal{L}|+|\mathcal{R}|)2^{\beta n(1-h(1/4))}\right)$ (Line 2). The $O^{*}$ here hides logarithmic factor in $|\mathcal{L}|+|\mathcal{R}|$ that grows polynomially in $n$.
Let us consider an iteration of the loop on line 2. Note that the values of $p$, $q$, $|\mathcal{Q}_{max}|$ and $\mu$ are now determined. The running time of Algorithm 2 on this specific iteration is
$O^{*}\left(q\cdot\left(|\mathcal{Q}_{max}|^{2}+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+(\frac{|\mathcal{Q}_{max}|^{4}}{pq}+2^{\beta nh(\max\{\mu,1/2-\mu\})})2^{\beta n(1-h(1/4))}\right)\right)\,.$
To bound the total running time of Algorithm 2, we can multiply the number of the iterations $\Theta^{*}(2^{\lambda n})$ by the upper bound on the slowest iteration.
Using Lemma 3.11, we bound the running time of Algorithm 2 as $O^{*}(2^{n/2}+2^{T(\alpha,\beta)\cdot n})$, where
$T(\alpha,\beta)=\frac{1}{2}\left(h(\gamma)-3\beta h(\gamma)+3\beta+2\lambda\right).$ (22)
Lemma 3.11.
The total running time of the Algorithm 2 is $O^{*}(2^{n/2}+2^{T(\alpha,\beta)\cdot n})$, where we define $T(\alpha,\beta)$ as follows:
$T(\alpha,\beta)=\frac{1}{2}\left(h(\gamma)-3\beta h(\gamma)+3\beta+2\lambda\right).$ (23)
Here, $\gamma=\gamma(\alpha)$ and $\lambda=\lambda(\alpha)$.
Proof.
As it was shown the total running time does not exceed
$O^{*}\left(2^{\lambda n}\cdot q\left(|\mathcal{Q}_{max}|^{2}+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+(\frac{|\mathcal{Q}_{max}|^{4}}{pq}+2^{\beta nh(\max\{\mu,1/2-\mu\})})2^{\beta n(1-h(1/4))}\right)\right)\;,$
where the values of $p$, $q$, $|\mathcal{Q}_{max}|$ and $\mu$ correspond to the slowest iteration of the loop on line 2.
Since $p$ and $q$ can vary only with polynomial factors, here by them we will denote the upper and lower bounds on their values, so let $\log p=\frac{(1-\varepsilon)}{2}\beta n$ and $\log q=\frac{1}{2}n-\lambda n-2\log|\mathcal{Q}_{max}|$.
Let us estimate the logarithm of the first summand inside the parenthesis:
$\log(2^{\lambda n}\cdot q\cdot|\mathcal{Q}_{max}|^{2})=\lambda n+\underbrace{\frac{1}{2}n-\lambda n-2\log|\mathcal{Q}_{max}|}_{\log q}+2\log|\mathcal{Q}_{max}|=\frac{1}{2}n.$
We estimate the last term of the sum (the fourth one) assuming that $\mu\geq\frac{1}{2}-\mu$, the other case can be shown similarly:
$\log\left(2^{\beta nh(\mu)+\beta n(1-h(1/4))}\right)=\beta nh(\mu)+\beta n-\beta nh(1/4)\leq 2\beta n\;.$
Recall that we choose $\beta<0.15$, then $2\beta n\leq 0.5n$.
The second summand is at most the third one, so it suffices to estimate the logarithm of the third summand. Recall that that $\log|\mathcal{Q}_{max}|\leq n\cdot\max\{\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)+\beta h\left(1/2-\mu\right)}{8},\;(1-\varepsilon)\beta\}$. If $\log|\mathcal{Q}_{max}|\leq(1-\varepsilon)\beta n$, then the third summand is dominated by $\frac{1}{2}(9\beta-2\beta h(\frac{1}{4}))$ and substituting $\beta=0.13$ yields value $\approx 0.48<0.5$. Therefore, this case is not considered further.
$\phantom{=}\log(2^{\lambda n}q\frac{|\mathcal{Q}_{max}|^{4}}{pq}2^{\beta n(1-h(1/4))})$
$=\lambda n+4\log|\mathcal{Q}_{max}|-\underbrace{\frac{1-\varepsilon}{2}\beta n}_{\log p}+\beta n-\beta nh(1/4)$
$=\lambda n+n\frac{h(\gamma)-3\beta h(\gamma)+3\beta-\varepsilon\beta+\beta h(\mu)+\beta h(1/2-\mu)-2\beta h(1/4)}{2}\;.$
Having $(-\varepsilon\beta)\leq 0$, $h(\mu)+h(1/2-\mu)\leq 2h(1/4)$, this expression can be dominated by
$\lambda n+n\frac{h(\gamma)-3\beta h(\gamma)+3\beta}{2}.$
So, the total running time of the algorithm is $O^{*}(2^{n/2}+2^{T(\alpha,\beta)\cdot n})$, where
$T(\alpha,\beta)=\frac{1}{2}\left(h(\gamma)-3\beta h(\gamma)+3\beta+2\lambda\right).$ (24)
We complete the analysis in Section 3.6.
3.5 Space Complexity
The space usage is dominated by the following parts.
  • Constructing $\mathcal{Q}_{1},\mathcal{Q}_{2},\mathcal{Q}_{3},\mathcal{Q}_{4}$ and $\mathcal{Q}_{1}^{\prime},\mathcal{Q}_{2}^{\prime},\mathcal{Q}_{3}^{\prime},\mathcal{Q}_{4}^{\prime}$ with $O^{*}(|\mathcal{Q}_{max}|)$ space (Line 2).
  • Finding $\mathcal{L},\mathcal{R}$ using Lemma 3.6 with $O^{*}\left(|\mathcal{Q}_{max}|+\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\max\{\mu,1/2-\mu\}|M|}\right)$ space (Line 2).
  • Solving $(\mathcal{L},\mathcal{R})$ using Lemma 2.2 with $O^{*}\left(|\mathcal{L}|+|\mathcal{R}|+2^{\beta n}\right)$ space (Line 2).
For a specific iteration of the loop on line 2, the space usage is at most the sum of the above expressions, where $|\mathcal{L}|,|\mathcal{R}|=O^{*}(\frac{|\mathcal{Q}_{max}|^{4}}{pq}+\binom{|M|}{\max\{\mu,1/2-\mu\}}|M|)$. The total space complexity of Algorithm 2 can be bounded by the space usage of an iteration, requiring the most space.
Therefore, it suffices to prove that $2^{\beta n}$, $\binom{|M|}{\max\{\mu,1/2-\mu\}|M|}$, $|\mathcal{Q}_{max}|$ and $\frac{|\mathcal{Q}_{max}|^{4}}{pq}$ are at most $O^{*}(2^{0.246n})$ on each iteration. By some simplifications (see Lemma 3.12) we can show that the total space usage of our algorithm is $O^{*}(2^{0.246n}+2^{S(\alpha,\beta)\cdot n})$, where
$S(\alpha,\beta)=\frac{1}{4}\left(3h(\gamma)-9\beta h(\gamma)+6\beta h(\frac{1}{4})+4\beta-2+4\lambda\right).$ (25)
Lemma 3.12.
The total space usage of Algorithm 2 is $O^{*}(2^{0.246n}+2^{S(\alpha,\beta)\cdot n})$, where $S(\alpha,\beta)$ is defined as follows:
$S(\alpha,\beta)=\frac{1}{4}\left(3h(\gamma)-9\beta h(\gamma)+6\beta h(\frac{1}{4})+4\beta-2+4\lambda\right).$ (26)
Here, we interpret $\gamma$ as $\gamma=\gamma(\alpha)$ and $\lambda$ as $\lambda(\alpha)$.
Proof.
As we have already discussed, it is sufficient to estimate the following values: $2^{\beta n}$, $\binom{|M|}{\max\{\mu,1/2-\mu\}|M|}$, $|\mathcal{Q}_{max}|$ and $\frac{|\mathcal{Q}_{max}|^{4}}{pq}$, for the iteration of the loop in line 2 that requires the most space. To do that, we estimate their logarithms. The first one is way smaller than $0.246n$, since $\beta<0.15$. The second one is equal to $\beta nh(\mu)\leq\beta n\leq 0.246n$.
Recall that that $\log|\mathcal{Q}_{max}|\leq n\cdot\max\{\frac{h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h(\mu)+\beta h\left(1/2-\mu\right)}{8},\;(1-\varepsilon)\beta\}$. If $\log|\mathcal{Q}_{max}|\leq(1-\varepsilon)\beta n$, then the third value is dominated by $\frac{1}{2}(11\beta-1)$ and substituting $\beta=0.13$ yields value $\approx 0.215<0.246$. Therefore, this case is not considered further.
$\log|\mathcal{Q}_{max}|=\frac{1}{8}\left(h(\gamma)-3\beta h(\gamma)+2\beta-2\varepsilon\beta+\beta h\left(\mu\right)+\beta h(\frac{1}{2}-\mu)\right)n\;.$
Having $(-2\varepsilon\beta)\leq 0$, $h(\mu)+h(1/2-\mu)\leq 2$ and $\beta<0.15$, we get
$\log|\mathcal{Q}_{max}|\leq\frac{1}{8}(1+\beta)n<0.246n.$
It is left to estimate $\log\frac{|\mathcal{Q}_{max}|^{4}}{pq}$.
$\log\frac{|\mathcal{Q}_{max}|^{4}}{pq}=4\log|\mathcal{Q}_{max}|-\underbrace{\frac{1-\varepsilon}{2}\beta n}_{p}-\underbrace{\frac{1}{2}n+2\log|\mathcal{Q}_{max}|+\lambda n}_{q}=$
$=\lambda n+n\frac{3h(\gamma)-9\beta h(\gamma)+3\beta h(\mu)+3\beta h(1/2-\mu)+4\beta-4\varepsilon\beta-2}{4}$
Having $(-4\varepsilon\beta)\leq 0$ and $h(\mu)+h(1/2-\mu)\leq 2h(1/4)$, we dominate this expression by
$\lambda n+\frac{3h(\gamma)-9\beta h(\gamma)+6\beta h(1/4)+4\beta-2}{4}\;.$
That concludes the proof. ∎
We complete the analysis in the following section.
3.6 Parameter Substitution
Lemma 3.13 (Maximum point for function of time complexity).
The function $T(\alpha,\beta)$ (see (22)) satisfies $T(\alpha,\beta)\leq T(0.5,0.13)$ when $\beta$ is defined as in (14).
Lemma 3.14 (Maximum point for function of space complexity).
The function $S(\alpha,\beta)$ (see (25)) satisfies $S(\alpha,\beta)\leq S(0.5,0.13)$ when $\beta$ is defined as in (14).
Both these functions contain a term $O(\frac{\log n}{n})$ that does not depend on the choice of $\alpha$ and $\beta$. In the proofs below, we omit this term.
Proof of Lemma 3.13.
Note that
$T(\alpha,\beta)$ $=\frac{1}{2}\left(h(\gamma)-3\beta h(\gamma)+3\beta+2(\alpha h\left(\frac{3\beta}{2\alpha}\right)+(1-\alpha)h\left(\frac{3\beta}{2(1-\alpha)}\right)-h(3\beta)))\right),$
$\frac{\partial^{2}T(\alpha,\beta)}{\partial\alpha^{2}}$ $=\frac{18\alpha^{2}\beta-2\alpha^{2}-18\alpha\beta-9\beta^{2}+2\alpha+6\beta}{\alpha(2\alpha-3\beta)(1-\alpha)(-2+3\beta+2\alpha)\ln(2)}.$
Consider two cases (the case $\alpha>0.5$ is symmetrical).
  1. 1.
    $0.45\leq\alpha\leq 0.5$.
    In this regime, we have $\beta=0.13$, the roots of the polynomials of the denominator and numerator are either $<0.2$ or $\geq 0.5$. By checking that $\frac{\partial T(\alpha,\beta)}{\partial\alpha}|_{a=\frac{1}{2}}=0$ and $\frac{\partial^{2}T(\alpha,\beta)}{\partial\alpha^{2}}|_{a=\frac{1}{2}}=0$, we verify that $a=\frac{1}{2}$ is indeed the maximum and have
    $T(\alpha,0.13)\leq T(0.5,0.13)\,.$
  2. 2.
    $\alpha<0.45$.
    In this case, $\beta=0$. Then all the roots of the polynomials of the denominator and numerator are either 00 or $1$. By checking that $\frac{\partial^{2}T(\alpha,\beta)}{\partial^{2}\alpha}|_{\alpha\in[0,0.45]}<0$ and $\frac{\partial T(\alpha,\beta)}{\partial\alpha}|_{\alpha\in[0,0.45]}>0$, we get that
    $T(\alpha,0)\leq T(0.45,0)\leq T(0.5,0.13)\,.$
Proof of Lemma 3.14.
Recall that
$S(\alpha,\beta)=\frac{1}{4}(3h(\gamma)-9\beta h(\gamma)+6\beta h\left(\frac{1}{4}\right)+4\beta-2+\\ 4(\alpha h\left(\frac{3\beta}{2\alpha}\right)+(1-\alpha)h\left(\frac{3\beta}{2(1-\alpha)}\right)-h(3\beta))).$
Consider two cases ($\alpha>0.5$ is symmetrical).
  1. 1.
    $0.45\leq\alpha\leq 0.5$. Then, $\beta=0.13$.
    $\frac{\partial^{2}S(\alpha,\beta)}{\partial\alpha^{2}}=-\frac{3(7\alpha^{2}\beta-\alpha^{2}-7\alpha\beta-3\beta^{2}+\alpha+2\beta)}{\alpha(\alpha-1)(2\alpha-3\beta)(3\beta+2\alpha-2)\ln(2)}$
    One can check that all the roots of the numerator and denominator are either $<0.2$ or $>0.5$. As $\frac{\partial^{2}S(\alpha,\beta)}{\partial\alpha^{2}}|_{\alpha=\frac{1}{2}}<0$, we conclude that $S$ is concave on the interval $[0.45,0.5]$. The only thing left to do is to verify that $\frac{\partial S(\alpha,\beta)}{\partial\alpha}|_{\alpha=\frac{1}{2}}=0$ which implies that $S(\alpha,0.13)\leq S(\frac{1}{2},0.13)$ if $0.45\leq\alpha\leq 0.5$.
  2. 2.
    $\alpha<0.45$. Then, $\beta=0$.
    $S(\alpha,0)=-\frac{1}{4}(3\alpha\log(\alpha)-3\alpha\log(1-\alpha)+2\log(2)+3\log(1-\alpha))$
    $\frac{\partial^{2}S(\alpha,0)}{\partial\alpha^{2}}=-\frac{3}{4\alpha(1-\alpha)\ln(2)}<0$
    Checking that $\frac{\partial S(\alpha,0)}{\partial\alpha}|_{\alpha\in[0,0.45]}>0$ we get that it is maximal at $\alpha=0.45$ and $S(\alpha,0)\leq S(0.45,0)\leq S(0.5,0.13)$. This concludes the proof.
Hence, the time and space complexity do not exceed $T(0.5,0.13)$ and $S(0.5,0.13)$, respectively (for any $\alpha$ and $\beta$ as in (14)).
Recall that $\lambda=h(3\beta)-\alpha h(\frac{3\beta}{2\alpha})-(1-\alpha)h(\frac{3\beta}{2(1-\alpha)})$, and $\gamma=\frac{\alpha-\frac{3}{2}\beta}{1-3\beta}$, see Equation (4).
By plugging $\alpha=0.5,\beta=0.13$, we get the following values for the parameters:
$\gamma$ $=0.5,$
$h(\gamma)$ $=1,$
$\lambda$ $=0,$
$T(0.5,0.13)$ $=\frac{1}{2}\cdot(1-3\cdot 0.13+3\cdot 0.13)=\frac{1}{2},$
$S(0.5,0.13)$ $\leq\frac{1}{4}\cdot(3-9\cdot 0.13+6\cdot 0.13\cdot 0.812+4\cdot 0.13-2)\leq 0.246.$
Combining Lemmas, we get that the algorithm runs in time $O^{*}\left(2^{0.5n}\right)$ and space $O^{*}\left(2^{0.246n}\right)$, which completes the proof of Theorem 1.1.
Acknowledgements
We would like to thank the anonymous reviewers who helped us to improve the quality of the text by finding a number of inconsistencies in the proofs as well as some missing references.